# DreamO: A Unified Framework for Image Customization

### v4

@@ 4.2. Qualitative Comparsion
To validate the performance of DreamO, we conduct comparisons with recent state-of-the-art methods across multiple subtasks. The visual comparison is shown in Fig. 5. The first part presents a comparison between DreamO and SOTA identity-customization methods i.e., PhotoMaker (Li et al., 2024), InstantID (Wang et al., 2024a), PuLID (Guo et al., 2024). The results demonstrate that DreamO can inject identity information with high fidelity across various scenes, while offering impressive flexibility for customization.
The second part compares DreamO with recent subject-customization methods, including single-task frameworks (i.e., MS-Diffusion (Wang et al., 2024b)) and unified generation frameworks (i.e., OmniGen (Xiao et al., 2024a), OminiControl (Tan et al., 2024)). The results demonstrate that DreamO achieves higher subject fidelity and better text consistency in both single-subject and multi-subbject scenarios.
@@
The last part presents the comparison between DreamO and recent style-customization methods, i.e., StyleShot (Gao et al., 2024), StyleAlign (Wu et al., 2021), InstantStyle (Wang et al., 2024c), DeaDiff (Qi et al., 2024), and CSGO (Xing et al., 2024). It can be observed that DreamO has weaker content intrusion, better text alignment, and higher style fidelity in the generated results.
@@ 4.3. Quantitative Comparison
In addition to qualitative comparison, we conduct quantitative comparisons for each task. We present the comparison of identity customization in Tab. 3, which is evaluated on Unsplash-50 (Gal et al., 2024). Here, we provide 9 prompts for each face. Following PuLID (Guo et al., 2024), the Face-Sim represents the ID cosine similarity, with ID embeddings extracted by CurricularFace (Huang et al., 2020). We also compute the CLIP cosine similarity between the generated result and the prompt to measure the text-following ability of different methods. As can be seen, our DreamO shows better face similarity and text-based customization ability.
Tab. 1 presents the quantitative comparison in single- and multi-subject customization. We use DreamBench (Ruiz et al., 2023) as the testset of single-subject customization. For multi-subject customization, we randomly select 20 pairs from DreamBench and provide 25 prompts for each. During testing, we generate four images with different seeds for each test sample. Here, we calculate the CLIP cosine similarity and Dino (Caron et al., 2021) cosine similarity between the generated result and the reference images as a measure of subject consistency. To improve accuracy, we remove the background of the generated results and then calculate similarity. Additionally, we employ CLIP cosine similarity between the text description and the generated result as a measure of content alignment. Tab. 1 shows that our method outperforms others in subject consistency while demonstrating strong text-following ability.
The quantitative comparison of the try-on is presented in Tab. 4. We select 300 reference garments from VITON-HD (Choi et al., 2021) encompassing various styles and colors as the test data. During testing, we provide 10 prompts for each cloth. The CLIP consine similarity between the generated result and reference cloth is employed to measure the try-on accuracy. Here, we only crop the cloth from the result to compute the CLIP-sim. The CLIP cosine similarity between the generated results and the prompt is employed to measure the text-following ability of different methods. The result in Tab. 4 shows the attractive performance of our DreamO. Although IMAGDressing (Shen et al., 2025) has higher CLIP-sim, it can only generate images with a white background (i.e., Fig. 5 ) with little text-following ability.
For style customization, we construct an evaluation dataset, containing $249$ style images and $24$ prompt. Each method generates $249\times 24$ style customization results. We use the pre-trained CSD (Somepalli et al., 2024) to extract the style feature of generated results and reference images, and calculate the cosine similarity between them as a measure of style consistency. Furthermore, we compute the CLIP (Radford et al., 2021) similarity between stylized results and text descriptions as the metric of content consistency. The results are shown in Tab. 2, which demonstrate that our method has better performance in style consistency and content consistency.
@@ 4.5. Ablation Study
Progressive training. To enable the model to better converge on all sub-tasks under complex data distributions and to rectify the impact of training data distribution on generation quality, we design a progressive training strategy. The effectiveness of this strategy is demonstrated in Fig. 8. One can see that directly training the model on all datasets leads to suboptimal convergence, particularly in complex tasks such as multi-subject consistency. Warming up on a smaller and easier-to-learn dataset (e.g., Subject200K (Tan et al., 2024)) before joint training improves convergence, but the generation quality is easily influenced by the training data distribution, deviating from the generation priors of Flux. By introducing an image quality tuning stage, the model can produce higher-quality generation results.

### v5

@@ 4.2. Qualitative Comparsion
To validate the performance of DreamO, we conduct comparisons with recent state-of-the-art methods across multiple subtasks. The visual comparison is shown in Fig. 5. The first part presents a comparison between DreamO and SOTA identity-customization methods i.e., PhotoMaker (photomaker), InstantID (instantid), PuLID (pulid). The results demonstrate that DreamO can inject identity information with high fidelity across various scenes, while offering impressive flexibility for customization.
The second part compares DreamO with recent subject-customization methods, including single-task frameworks (i.e., MS-Diffusion (ms_diffusion)) and unified generation frameworks (i.e., OmniGen (omnigen), OminiControl (ominicontrol)). The results demonstrate that DreamO achieves higher subject fidelity and better text consistency in both single-subject and multi-subbject scenarios.
@@
The last part presents the comparison between DreamO and recent style-customization methods, i.e., StyleShot (styleshot), StyleAlign (stylealign), InstantStyle (wang2024instantstyle), DeaDiff (deadiff), and CSGO (csgo). It can be observed that DreamO has weaker content intrusion, better text alignment, and higher style fidelity in the generated results.
@@ 4.3. Quantitative Comparison
In addition to qualitative comparison, we conduct quantitative comparisons for each task. We present the comparison of identity customization in Tab. 3, which is evaluated on Unsplash-50 (usp). Here, we provide 9 prompts for each face. Following PuLID (pulid), the Face-Sim represents the ID cosine similarity, with ID embeddings extracted by CurricularFace (cface). We also compute the CLIP cosine similarity between the generated result and the prompt to measure the text-following ability of different methods. As can be seen, our DreamO shows better face similarity and text-based customization ability.
Tab. 1 presents the quantitative comparison in single- and multi-subject customization. We use DreamBench (dreambooth) as the testset of single-subject customization. For multi-subject customization, we randomly select 20 pairs from DreamBench and provide 25 prompts for each. During testing, we generate four images with different seeds for each test sample. Here, we calculate the CLIP cosine similarity and Dino (dino) cosine similarity between the generated result and the reference images as a measure of subject consistency. To improve accuracy, we remove the background of the generated results and then calculate similarity. Additionally, we employ CLIP cosine similarity between the text description and the generated result as a measure of content alignment. Tab. 1 shows that our method outperforms others in subject consistency while demonstrating strong text-following ability.
The quantitative comparison of the try-on is presented in Tab. 4. We select 300 reference garments from VITON-HD (viton) encompassing various styles and colors as the test data. During testing, we provide 10 prompts for each cloth. The CLIP consine similarity between the generated result and reference cloth is employed to measure the try-on accuracy. Here, we only crop the cloth from the result to compute the CLIP-sim. The CLIP cosine similarity between the generated results and the prompt is employed to measure the text-following ability of different methods. The result in Tab. 4 shows the attractive performance of our DreamO. Although IMAGDressing (imdress) has higher CLIP-sim, it can only generate images with a white background (i.e., Fig. 5 ) with little text-following ability.
For style customization, we construct an evaluation dataset, containing $249$ style images and $24$ prompt. Each method generates $249\times 24$ style customization results. We use the pre-trained CSD (csd) to extract the style feature of generated results and reference images, and calculate the cosine similarity between them as a measure of style consistency. Furthermore, we compute the CLIP (clip) similarity between stylized results and text descriptions as the metric of content consistency. The results are shown in Tab. 2, which demonstrate that our method has better performance in style consistency and content consistency.
@@ 4.5. Ablation Study
Progressive training. To enable the model to better converge on all sub-tasks under complex data distributions and to rectify the impact of training data distribution on generation quality, we design a progressive training strategy. The effectiveness of this strategy is demonstrated in Fig. 8. One can see that directly training the model on all datasets leads to suboptimal convergence, particularly in complex tasks such as multi-subject consistency. Warming up on a smaller and easier-to-learn dataset (e.g., Subject200K (ominicontrol)) before joint training improves convergence, but the generation quality is easily influenced by the training data distribution, deviating from the generation priors of Flux. By introducing an image quality tuning stage, the model can produce higher-quality generation results.
@@ 5. Conclusion
Figure 10. The capability of our proposed DreamO in identity-driven image customization.
Figure 11. The capability of our proposed DreamO in subject-driven image customization.
Figure 12. The capability of our proposed DreamO in try-on image customization.
Figure 13. The capability of our proposed DreamO in style-driven image customization.